OpenStudy (anonymous):

A car is travelling at 72km/h and the brakes are fully applied, producing a constant deceleration of 12m/s^2. a) Verify that the velocity function v(t)=-12t+20, where t = seconds, gives this deceleration and initial velocity. b) How long does it take for the car to come to a complete stop?

OpenStudy (tylerd):

so first convert 72km an hour to m/s

OpenStudy (tylerd):

we get 20m/s

OpenStudy (tylerd):

the car comes to a stop when its velocity is 0

OpenStudy (tylerd):

-12t+20 = 0

OpenStudy (tylerd):

solve for t

OpenStudy (anonymous):

t would equal to 5/3? @TylerD

OpenStudy (tylerd):

20/12 = 5/3 yes

OpenStudy (tylerd):

to verify the velocity function, the derivative of it is -12, which is acceleration.

OpenStudy (tylerd):

and because its in m/s^2, we had to convert the km/hr to m/s

OpenStudy (tylerd):

you can also get the position function, by integrating. p(t)=-6t^2+20t

OpenStudy (tylerd):

but no need, just saying..

OpenStudy (anonymous):

oh okay! thanks a lot! If you don't mind, can you explain what integrating? I'm curious

OpenStudy (triciaal):

@TylerD you are thinking, good! but you have an error you integrate the velocity to get the distance same as differentiate the distance to get the velocity @esam2 basically integration and differentiation are opposites like multiplication and division

OpenStudy (triciaal):

when you differentiate you are finding the rate of change in essence the slope as the change in the independent variable approaches zero

OpenStudy (triciaal):

@TylerD sorry, I read the post without the period and you did not say what you were integrating. I thought you were integrating the Position function P(t). don't forget your C when you integrate